# A Data-Driven Approach for Topology Correction in Low Voltage Networks with DERs

### v1

@@ II-C Phase Label Identification
II-D Time-based Data Selection
Compared to traditional LVDNs, the voltage magnitude measurements in modern LVDNs may exhibit higher values due to the integration of PV systems. The injection of residual power from PV-equipped users can lead to a scenario where the user voltage exceeds the surrounding voltages. Within the same distribution network area, environmental factors such as light intensity, temperature, and humidity remain relatively uniform, leading to similar PV generation patterns among users. This effect is particularly pronounced during summer when high-power PV systems are connected. Consequently, users with identical PV capacities but connected to different phases may experience analogous voltage magnitude fluctuations. This phenomenon introduces challenges to the identification steps based on correlation analysis, as the voltage magnitude variations induced by PV generation may obscure the underlying phase-dependent characteristics.
@@ II-D Time-based Data Selection
III Case study
In this section, the feasibility and robustness of the proposed topology correction framework are validated using modified real-world LVDNs and SM data collected from customers in the Netherlands. The resolution of the dataset is 15 minutes, resulting in 96 data points per day. The base three-phase voltage level is 0.4 kV, and the loads of three-phase customers are non-uniformly distributed across Phases A, B, and C. The time thresholds $t_{1}$ and $t_{2}$ are set to 20 and 88. The voltage magnitude profiles are generated using a power flow model, being solved by using the Power Grid Model package[32].
III-A Switch State Analysis
TABLE II: Accuracy (%) comparison under multiple scenarios
@@ III-A Switch State Analysis
The switch bar is connected to three and five transformers, respectively, with all transformers assumed to be connected to the same MV network, which introduces similar low-frequency variations in customer voltage profiles. Each feeder connects to 49 customers. Using the voltage datasets selected through the time-based strategy described in Section II-D, the performance of three models is evaluated under varying numbers of available SMs at each feeder, including RF, Support Vector Machine (SVM), and Gaussian Naive Bayes (GNB).
@@
Table II demonstrates that the accuracy of all methods improves as the number of available SM increases. SVM achieves high accuracy (i.e., above 98% for 5 switches scenarios and 88% for 3 switches scenarios) once the number of houses reaches 10, while RF reaches the same level only when all SM data are available. This suggests that SVM is more efficient in learning from limited data, likely due to its ability to find optimal decision boundaries even with fewer samples. For the two SM data available scenarios, the accuracy of RF is about 5% higher than that of SVM in two scenarios (i.e., 3 and 5 switches in the switch bar). Moreover, GNB underperforms compared to RF and SVM, with its accuracy lagging behind when the number of houses is low (lower than 90% and 95% under all the SMs data in two cases, respectively).
Figure 8: PCC in first row and MFP in second row among voltage datasets: a) with single-phase customers and b) with three-phase customers
III-B User-feeder Identification Analysis
The influence of network complexity, phase configuration, and per-node connection on classification accuracy was first analyzed. The per-node connection is the number of households connected to a point, e.g., a multi-unit building. Given datasets with SM errors but without available recordings of user-feeder connections, Table II presents the accuracy comparison under five modified configurations of LVDNs: a network with two complex-structured feeders (CN1), a network with a simple-structured and one complex-structured feeder (CN2), a network with two simple-structured feeders (SN), network with longer simple-structured feeders (SNL), and network with two simple-structured feeders with more branches (SNB). Another complex-structured feeder or simple-structured feeder is added to the above five LVDNs to construct LVDNs with three feeders (as shown in the fifth column). The main difference between a complex network and a simple network is the length of feeders and the number of branches. The topology of CN2 and SNB are shown in Fig.6.
III-B1 Identification Without Recordings
TABLE III: Accuracy (%) comparison under multiple scenarios
@@ III-B1 Identification Without Recordings
The PCC and MFP among customers in CN1 with multiple connections per node are depicted in Fig. 8. As illustrated in Fig. 8(a), expression in (7) reduces the voltage correlation between users connected to different feeders. This step helps distinguish between different connections more clearly. However, the complexity of the voltage data correlation shown in Fig. 8(b)—reflected in the similarity of the correlation distribution — makes it extremely challenging to simultaneously identify both the phase label and the user-feeder connection using only voltage data. A balanced LVDN represents a scenario where all users are connected to the same phase (i.e., a single-phase LVDN), whereas an unbalanced LVDN corresponds to a three-phase configuration.
TABLE IV: Average accuracy under three-phase datasets and multiple connections at nodes with SM errors
@@
Table III summarizes the accuracy of user-feeder connection identification across five LVDN scenarios. In general, the proposed approach demonstrates higher robustness to variations in the number of feeders connected to the same transformer and per-node connections but exhibits lower robustness in three-phase configurations. Specifically, accuracy drops from 100% to 96.43% in CN1 with three feeders and multiple connections per node, and from 94.56% to 55.56% in SN with two feeders and multiple connections per node, highlighting the impact of complex multi-phase settings. In simpler LVDNs, per-node connections significantly affect accuracy. For instance, in SNL with two feeders in a single-phase setting, accuracy declines from 96.3% to 82.22%, while in SNL with two feeders in a three-phase setting, it drops from 59.26% to 46.67%. In contrast, for complex LVDNs, accuracy remains above 95% regardless of single or multiple per-node connections, indicating greater resilience in more intricate network structures. Moreover, accuracy in complex LVDNs is slightly higher than in simpler LVDNs, primarily due to feeder length differences. Simple LVDNs, which are modified from urban areas with high population density, have houses located close to each other (i.e., smaller distances between users with different connections). This proximity reduces the difference in the voltage variations between neighbouring houses, leading to greater identification accuracy. However, the length of the feeder and the number of branches have a minor effect on the accuracy, which is verified in the three simple LVDNs (i.e., SN, SNL, and SNB).
@@
III-B2 Identification With Recordings
Given datasets with varying levels of unknown user-feeder connections, Table IV compares the average accuracy of 20 simulations at each scenario (i.e., each unknown level). The results provide insights into the robustness and effectiveness of both methods across different network configurations. In general, as the percentage of unknown user-feeder connections increases from 10% to 50%, a gradual decline in accuracy is observed across all methods and network configurations. However, both the proposed MFP-based method and KNN (with varying $k$ values) maintain relatively high performance, with accuracy exceeding 91% even at 30% unknown connections.
Figure 9: Accuracy under incomplete dataset: a) modified KNN with 20% unknown labels, b) traditional KNN with 20% unknown labels, c) modified KNN with 50% unknown labels, b) traditional KNN with 50% unknown labels.
Table IV indicates that the traditional KNN algorithm consistently outperforms the proposed modified KNN approach when the parameter $k$ is less than 8. However, the modified KNN demonstrates improved robustness as the proportion of missing data increases, particularly at higher unknown rates and a larger $k$. Specifically, the modified KNN method outperforms traditional KNN at higher levels of unknown information (e.g., 50%), achieving accuracy levels exceeding 87%. This suggests the proposed method may better preserve identification accuracy under limited accurate recordings from DSOs. In the simple network with branches (SNB), the performance gap between the two methods becomes less pronounced. Traditional KNN shows a slight advantage at intermediate unknown rates (e.g., 15%–20%) and smaller $k$ values, whereas the modified KNN achieves comparable or superior performance at higher levels of missing data and larger $k$ values. Fig. 9 depicts the accuracy of the approaches under datasets with various missing data percentages. The accuracy declines slightly about 5% as the percentage of missing data increases, even at 70% missing data. This demonstrates the strong robustness of the proposed method against missing data.
TABLE V: Average purity (%) under datasets form feeders with different components
@@ III-B2 Identification With Recordings
III-C Phase Identification Analysis
The influence of feeder complexity, environmental conditions, and load distribution on identification accuracy is analyzed, with results summarized in Table V and VI. CC_PV: Represents a complex-structured feeder from CN1, with photovoltaics (PV) installed at each house. SC_PV: Corresponds to a simple-structured feeder from SN, also with PV at each house. CC and SC: Represent feeders from CN1 and SN with PV, but with datasets preprocessed using the method described in Section II-D. These four feeders serve single-phase users only.CC_3_NU and SC_3_NU: Denote complex and simple feeders equipped with three-phase meters, where the load is non-uniformly distributed among the three phases. “+20” indicates an additional 20 customers connected to the feeder. “Long feeder” refers to an extended feeder length compared to the original configuration. Accuracy and clustering purity are used to evaluate the performance of phase identification, as formulated in expression (15).
[…]
TABLE VI: Average accuracy (%) under datasets form feeders with different components.
@@ III-C Phase Identification Analysis
Overall, Table V indicate consistently higher accuracy in winter datasets compared to summer datasets. This discrepancy is primarily due to voltage variations caused by household photovoltaic (PV) systems, which negatively impact phase identification performance. In the first four cases, where only single-phase or balanced three-phase customers are considered, the influence of PV-induced voltage fluctuations is evident. For instance, in the summer dataset, the CC_PV feeder (a complex feeder from CN1 without PV filtering) achieves 88.89% accuracy with a single connection per node but drops to 72.08% with multiple connections per node. However, when PV-induced variations are filtered out—as in the CC and SC cases—accuracy consistently exceeds 95%, highlighting the critical role of PV filtering in improving phase identification.
@@
IV Conclusion
In this paper, we proposed a practical topology correction framework for LVDNs, incorporating a time-based smart meter selection strategy to mitigate the impact of PV systems and a balanced dataset construction to improve training effectiveness. The experimental results demonstrate that the proposed framework is highly robust, maintaining high accuracy even with incomplete datasets, regardless of random missing data, data dimensionality, or topological complexity. While the correlation-based user and phase identification shows strong resilience to data loss, the switch state identification component, based on supervised learning, is more sensitive to missing data. Moreover, the framework’s performance is influenced by seasonal variations and voltage fluctuations caused by DERs, highlighting the importance of time-based data selection in ensuring reliable topology identification.

### v2

@@ Introduction
Time-based Data Selection
Compared to traditional LVDNs, the voltage magnitude measurements in modern LVDNs may exhibit higher values due to the integration of PV systems. Within the same distribution network area, environmental factors such as light intensity, temperature, and humidity remain relatively uniform, leading to similar PV generation patterns among users. The voltage magnitudes from the same LVDN are depicted in Fig. 6. This effect is particularly pronounced during summer when high-power PV systems are connected (i.e., the voltage magnitude distribution in Fig. 6(b)). Consequently, users with identical PV capacities but connected to different phases may experience analogous voltage magnitude fluctuations. This phenomenon introduces challenges to the identification steps based on correlation analysis, as the voltage magnitude variations induced by PV generation may obscure the underlying phase-dependent characteristics.
@@
Case Study
In this section, the feasibility and robustness of the proposed topology correction framework are validated using modified real-world LVDNs and SM data collected from customers in the Netherlands. The code used in this paper is available online  [^1] https://github.com/distributionnetworksTUDelft/Data-driven-approach-for-topology-correction-in-LVDNs-. The resolution of the dataset is 15 minutes, resulting in 96 data points per day. The base three-phase voltage level is 0.4 kV, and the loads of three-phase customers are non-uniformly distributed across Phases A, B, and C. The time thresholds $t_{1}$ and $t_{2}$ are set to 20 and 88. The voltage magnitude profiles are generated using a power flow model, being solved by using the Power Grid Model package [PowerGridModel].
Switch State Analysis
Table 2: Accuracy (%) comparison under multiple scenarios
@@
The switch bar is connected to three and five transformers, respectively, with all transformers assumed to be connected to the same MV network, which introduces similar low-frequency variations in customer voltage profiles. Each feeder connects to 49 customers. Using the voltage datasets selected through the time-based strategy described in Section 2.4, the performance of three models is evaluated under varying numbers of available SMs at each feeder, including RF, Support Vector Machine (SVM), and Gaussian Naive Bayes (GNB).
Table 2 demonstrates that the accuracy of all methods improves as the number of available SM increases. SVM achieves high accuracy (i.e., above 98% for 5 switches scenarios and 88% for 3 switches scenarios) once the number of houses reaches 10, while RF reaches the same level only when all SM data are available. This suggests that SVM is more efficient in learning from limited data, likely due to its ability to find optimal decision boundaries even with fewer samples. For the two SM data available scenarios, the accuracy of RF is about 5% higher than that of SVM in two scenarios (i.e., 3 and 5 switches in the switch bar). Moreover, GNB underperforms compared to RF and SVM, with its accuracy lagging behind when the number of houses is low (lower than 90% and 95% under all the SMs data in two cases, respectively).
@@
Table 3: Accuracy (%) comparison under multiple scenarios
@@
User-feeder Identification Analysis
The influence of network complexity, phase configuration, and per-node connection on classification accuracy was first analysed. The per-node connection is the number of households connected to a point, e.g., a multi-unit building. Given datasets with SM errors but without available recordings of user-feeder connections, Table 3 presents the accuracy comparison under five modified configurations of LVDNs: a network with two complex-structured feeders (CN1), a network with a simple-structured and one complex-structured feeder (CN2), a network with two simple-structured feeders (SN), network with longer simple-structured feeders (SNL), and a network with two simple-structured feeders with more branches (SNB). Another complex-structured feeder or simple-structured feeder is added to the above five LVDNs to construct LVDNs with three feeders (as shown in the fifth column). The main difference between a complex network and a simple network is the length of feeders and the number of branches. The topology of CN2 and SNB are shown in Fig. 8.
Figure 9: PCC in first row and FPCC in second row among voltage datasets: a) with single-phase customers and b) with three-phase customers
Table 4: Average accuracy under three-phase datasets and multiple connections at nodes with SM errors
@@
Identification Without Recordings
The PCC and MFP among customers in CN1 with multiple connections per node are depicted in Fig. 9. As illustrated in Fig. 9(a), the expression in (7) reduces the voltage correlation between users connected to different feeders. This step helps distinguish between different connections more clearly. However, the complexity of the voltage data correlation shown in Fig. 9(b)—reflected in the similarity of the correlation distribution — makes it extremely challenging to simultaneously identify both the phase label and the user-feeder connection using only voltage data. A balanced LVDN represents a scenario where all users are connected to the same phase (i.e., a single-phase LVDN), whereas an unbalanced LVDN corresponds to a three-phase configuration.
Table 3 summarizes the accuracy of user-feeder connection identification across five LVDN scenarios. In general, the proposed approach demonstrates higher robustness to variations in the number of feeders connected to the same transformer and per-node connections but exhibits lower robustness in three-phase configurations. Specifically, accuracy drops from 100% to 96.43% in CN1 with three feeders and multiple connections per node, and from 94.56% to 55.56% in SN with two feeders and multiple connections per node, highlighting the impact of complex multi-phase settings. In simpler LVDNs, per-node connections significantly affect accuracy. For instance, in SNL with two feeders in a single-phase setting, accuracy declines from 96.3% to 82.22%, while in SNL with two feeders in a three-phase setting, it drops from 59.26% to 46.67%. In contrast, for complex LVDNs, accuracy remains above 95% regardless of single or multiple per-node connections, indicating greater resilience in more intricate network structures. Moreover, accuracy in complex LVDNs is slightly higher than in simpler LVDNs, primarily due to feeder length differences. Simple LVDNs, which are modified from urban areas with high population density, have houses located close to each other (i.e., smaller distances between users with different connections). This proximity reduces the difference in the voltage variations between neighbouring houses, leading to greater identification accuracy. However, the length of the feeder and the number of branches have a minor effect on the accuracy, which is verified in the three simple LVDNs (i.e., SN, SNL, and SNB).
@@
Identification With Recordings
Given datasets with varying levels of unknown user-feeder connections, Table 4 compares the average accuracy of 20 simulations at each scenario (i.e., each unknown level). The results provide insights into the robustness and effectiveness of both methods across different network configurations. In general, as the percentage of unknown user-feeder connections increases from 10% to 50%, a gradual decline in accuracy is observed across all methods and network configurations. However, both the proposed MFP-based method and KNN (with varying $k$ values) maintain relatively high performance, with accuracy exceeding 91% even at 30% unknown connections.
Table 4 indicates that the traditional KNN algorithm consistently outperforms the proposed modified KNN approach when the parameter $k$ is less than 8. However, the modified KNN demonstrates improved robustness as the proportion of missing data increases, particularly at higher unknown rates and a larger $k$. Specifically, the modified KNN method outperforms traditional KNN at higher levels of unknown information (e.g., 50%), achieving accuracy levels exceeding 87%. This suggests the proposed method may better preserve identification accuracy under limited accurate recordings from DSOs. In the simple network with branches (SNB), the performance gap between the two methods becomes less pronounced. Traditional KNN shows a slight advantage at intermediate unknown rates (e.g., 15%–20%) and smaller $k$ values, whereas the modified KNN achieves comparable or superior performance at higher levels of missing data and larger $k$ values. Fig. 10 depicts the accuracy of the approaches under datasets with various missing data percentages. The accuracy declines slightly by about 5% as the percentage of missing data increases, even at 70% missing data. This demonstrates the strong robustness of the proposed method against missing data.
Figure 10: Accuracy under incomplete dataset: a) modified KNN with 20% unknown labels, b) traditional KNN with 20% unknown label, c) modified KNN with 50% unknown labels, d) traditional KNN with 50% unknown labels.
Phase Identification Analysis
The influence of feeder complexity, environmental conditions, and load distribution on identification accuracy is analyzed, with results summarized in Table 5 and 6. CC_PV: Represents a complex-structured feeder from CN1, with photovoltaics (PV) installed at each house. SC_PV: Corresponds to a simple-structured feeder from SN, also with PV at each house. CC and SC: Represent feeders from CN1 and SN with PV, but with datasets preprocessed using the method described in Section 2.4. These four feeders serve single-phase users only.CC_3_NU and SC_3_NU: Denote complex and simple feeders equipped with three-phase meters, where the load is non-uniformly distributed among the three phases. “+20” indicates an additional 20 customers connected to the feeder. “Long feeder” refers to an extended feeder length compared to the original configuration. Accuracy and clustering purity are used to evaluate the performance of phase identification, as formulated in the expression (15).
$P_{pu}=\frac{1}{N}\sum_{\mathcal{L}_{k}\in{\mathcal{L}}}\underset{i}{\operatorname{max}}|\mathcal{L}_{k}\cap\mathcal{L}_{i}|\$(15)
Table 5: Average purity (%) under datasets form feeders with different components
@@
Table 6: Average accuracy (%) under datasets form feeders with different components.
@@
Overall, Table 5 indicates consistently higher accuracy in winter datasets compared to summer datasets. This discrepancy is primarily due to voltage variations caused by household photovoltaic (PV) systems, which negatively impact phase identification performance. In the first four cases, where only single-phase or balanced three-phase customers are considered, the influence of PV-induced voltage fluctuations is evident. For instance, in the summer dataset, the CC_PV feeder (a complex feeder from CN1 without PV filtering) achieves 88.89% accuracy with a single connection per node but drops to 72.08% with multiple connections per node. However, when PV-induced variations are filtered out—as in the CC and SC cases—accuracy consistently exceeds 95%, highlighting the critical role of PV filtering in improving phase identification.
@@
Conclusion
In this paper, we proposed a practical topology correction framework for LVDNs, incorporating a time-based smart meter selection strategy to mitigate the impact of PV systems and a balanced dataset construction to improve training effectiveness. The experimental results demonstrate that the proposed framework is highly robust, maintaining high accuracy even with incomplete datasets, regardless of random missing data, data dimensionality, or topological complexity. Specifically, the correlation-based user and phase identification shows strong resilience to data loss. However, the switch state identification component, based on supervised learning, is more sensitive to missing and asynchronous data. Moreover, the framework’s performance is influenced by seasonal variations and voltage fluctuations caused by PVs, highlighting the importance of time-based data selection in ensuring reliable topology identification. Future extensions of this work will account for the influence of various flexible assets to enhance the scalability and precision of the framework.
[…]